The Eiffel tower, which is made of steel (coefficient of linear expansion 13 x 10^-6/ degree C.), is 320 m high at 20(Degrees C). the change in height of the tower between the winter months when the average temperature is -20C and the summer months when the average temperature is 35C is ?

a. 42 cm

b. 35 cm

c. 23cm

d. 15 cm

e. 11cm
Answer #1

Change in height = L*coefficient of linear expansion *Delta T

= 320*13*10^-6*(35 +20) = 0.2288 m = 23 cm

Answer : Option C
Answer #2

change = 320*13*10^-6*(20+35)
=23 cm